# Do massive neutrino states really exist?

### v1

@@ Abstract
In neutrino physics, it is conventional to work with massive states corresponding to defined neutrino masses. However, in weak interaction processes, only flavor states participate, and the propagation can be described by an off-diagonal mass matrix. In this work, we operate directly with the non-diagonal mass matrix and employ a quantum field theory approach to describe neutrino oscillations within the plane-wave approximation, circumventing the use of wave packets. Within this framework, numerical estimates for the amplitude of a local process involving lepton interactions via virtual neutrino exchange yield values of order $10^{-53}$. For macroscopic processes with neutrino exchange, application of the matrix propagator reproduces the standard quantum mechanical neutrino oscillation probability formula while preserving relativistic invariance.
@@ Introduction
Typically, the neutrino which is propagating for the long distances, is considered as a real particle and thus must exist in a mass eigenstate. This is achieved through the diagonalization of the mass matrix entering the respective Lagrangian. However, within the modern framework, it is treated as a virtual particle. This then enables the computation of these propagators without having to diagonalize the mass matrix.
In this work, we extend the QFT approach to processes with lepton number violation, analyzing a two-flavor model of Dirac neutrinos. Unlike the work of Kovalenko and $\check{\text{S}}$imkovic [3] , who derived diagonal propagators for massive neutrinos, we consider flavor-neutrino propagation which are described by a non-diagonal matrix propagator. A similar methodology was employed by M. Dvornikov [4] in analyzing oscillations in matter.
Next, we consider the mass matrix in the Chau-Keung parameterization of the PMNS matrix via Euler angles. We discuss how parameterization can affect the normal and inverted neutrino mass hierarchies.
We demonstrate that local weak processes (e.g., $l^{-}_{\alpha}l^{+}_{\beta}\to W^{+}W^{-}$) exhibit oscillation-sensitive contributions suppressed by $\mathcal{O}(10^{-79})$. For macroscopic diagrams with two-flavor transitions, the derived amplitude reproduces the canonical oscillation dependence
$P_{osc}\propto\sin^{2}\left(\cfrac{\Delta m^{2}L}{4E_{\text{avg}}}\right),$
where the oscillation phase is directly linked to neutrino masses. This formalism bridges QFT foundations and experimental interpretation while simultaneously confirming consistency with quantum mechanical results.
@@ 1 Dirac Neutrino Propagator
The propagator in momentum space, analogous to [1], is defined as $\mathbb{K}^{-1}(p)\equiv\mathbb{D}(p)$ :
@@ 1.1 Two-Flavor Propagator
We consider the specific case of interaction between two flavors, electron ($e$) and muon ($\mu$), with the mass matrix:
@@ 2 Three flavor mass matrix
Analyzing the propagator for three neutrino flavors in its off-diagonal form presents several complications. However, it is useful to trace the connection between the mass matrix and the mixing angles for three flavors. When mass matrix have only real elements, we have 6 independent variables that can be expressed through 6 parameters: 3 masses neutrino and 3 mixing angles. But we have a Hermitian mass matrix with nine different variables and (2.2) is expressed through the only seven independent parameters: 3 masses, 3 mixing angles and CP-phase $\delta$. We expect there are two relation between mass matrix elements. In the general case, the mass matrix is Hermitian:
@@
2.1 Chau-Keung parametrization
Standard Chau-Keung paramtrization of Pontecorvo-Maki-Nakagawa-Sakata (PMNS) matrix $V_{\text{PMNS}}$ has three Euler angles $\theta_{12},\theta_{23},\theta_{13}$ and Dirac CP-phase $\delta$:
@@ 2.2 Numerical estimates
For numerical estimates, we use the standard Chau-Keung parametrization, in which the angles and the CP-violating phase have been measured experimentally. We have 2 different situation for Inverted and Normal Ordering mass states.
@@
$\theta_{12}\,/^{\circ}$ $33.68^{+0.73}_{-0.70}$ $33.68^{+0.73}_{-0.70}$
$\theta_{23}\,/^{\circ}$ $43.3^{+1.0}_{-0.8}$ $47.9^{+0.7}_{-0.9}$
@@
$\cfrac{\Delta m^{2}_{31}}{10^{-3}\,\text{eV}^{2}}$ $+2.513^{+0.021}_{-0.019}$ $-2.409^{+0.020}_{-0.020}$
$\cfrac{\Delta m^{2}_{32}}{10^{-3}\,\text{eV}^{2}}$ $+2.438^{+0.021}_{-0.019}$ $-2.484^{+0.020}_{-0.020}$
Table 1: NuFIT 6.0 (2024): Oscillation parameters IC24 [10]
Figure 1: Mandelstam plane analog
One of the mixing angles is small compared to the others. However, how can an increase in this angle affect the values of the mass matrix? What happens to the diagonal elements of this matrix? Will they increase monotonically or not, and when might this behavior change?
Normal hierarchy ($\mathbb{M}_{\text{NH}}\times 10^{-2}\text{ eV}$)
$\mathbb{M}_{\text{NH}}(m_{1}=0)\simeq\begin{pmatrix}0.371&-0.119+0.254i&-0.702+0.269i\\ -0.119-0.254i&2.676&2.153\,-0.032i\\ -0.702-0.269i&2.153\,+0.032i&2.831\\ \end{pmatrix}$ (2.6)
Inverse hierarchy ($\mathbb{M}_{\text{IH}}\times 10^{-2}\text{ eV}$):
$\mathbb{M}_{\text{IH}}(m_{3}=0)\simeq\begin{pmatrix}4.822&-0.015-0.539i&-0.06-0.487i\\ -0.015+0.539i&2.29&-2.413-0.005i\\ -0.06+0.487i&-2.413+0.005i&2.781\\ \end{pmatrix}$ (2.7)
Substituting the values for the angles and masses into the expression for the mass, it was observed that for the normal hierarchy (NH), the values of the diagonal elements of the mass matrix form a monotonically increasing sequence. However, upon increasing the mixing angle $\theta_{13}$ and exceeding $\theta_{13}\simeq 19.8^{\circ}$, the sequence ceases to be monotonically increasing. The same occurs when changing the masses: if the absolute difference between $m_{1}$ and $m_{2}$ is larger than their difference from $m_{3}$, the sequence of diagonal elements of the mass matrix also ceases to be monotonically increasing.
For the inverted hierarchy (IH), the sequence of elements on the main diagonal of the mass matrix was initially neither monotonically increasing nor decreasing. However, upon reaching an angle $\theta_{13}\simeq 36.5^{\circ}$ or larger, the values on the main diagonal formed a monotonically increasing sequence:
(3.186, 3.188, $(3.186,\ 3.188,\ 3.517)\cdot 10^{-2}.$ (2.8)
Substituting known mixing angles and masses in (2.2), we obtain the dependence plots of neutrino mass matrix elements on the lightest neutrino mass for both mass hierarchies (AppendixA).
@@
$\begin{split}&m_{3}=0\times 10^{-2}\text{ eV}\ \sim\ \Re(m_{e\mu})=-15\cdot 10^{-3},\\ &m_{3}=2\times 10^{-2}\text{ eV}\ \sim\ \Re(m_{e\mu})=-4\cdot 10^{-3},\\ &m_{3}=4\times 10^{-2}\text{ eV}\ \sim\ \Re(m_{e\mu})=-6\cdot 10^{-5}.\end{split}$ (2.9)
@@
Figure 2: Plots of Inverse (left) and Normal (right) Hierarchy as a function of lightest neutrino mass (IH:$m_{3}$, NH:$m_{1}$).
@@
All non-diagonal elements of neutrino mass matrix exhibit power-law decay towards zero without sign inversion. However, $\Re(m_{e\mu})$ displays distinct behavior in both mass hierarchies. For the Normal Hierarchy (NH), we observe an initial local minimum followed by asymptotic convergence to zero. In the Inverted Hierarchy (IH), $\Re(m_{e\mu})$ undergoes sign reversal, reaches a global maximum beyond cosmological bounds on $m_{3}$, and subsequently approaches zero from above. This pronounced peak, while cosmologically unphysical, presents a striking feature. The inverse hierarchy is related to some kind of fine-tunning for matrix. This may be an additional argument in choosing the normal hierarchy (NH).
@@ 3 Weak Interaction Process
where $\mathcal{P}_{L}=(1-\gamma_{5})/2$ are the left-handed projectors, satisfying $\mathcal{P}_{L}\mathcal{P}_{R}=0$. In Feynman diagrams, vertices include $\gamma^{\mu}\mathcal{P}_{L}$. Consequently, in the amplitude squared, terms with constants between vertices vanish due to:
@@ 3.1 Process $l_{\alpha}^{-}l_{\beta}^{+}\to W^{+}W^{-}$
We consider the process of two leptons interacting to form two W-bosons and evaluate the contribution of oscillations to this local process.
Figure 3: Diagrams of $l_{\alpha}^{-}l_{\beta}^{+}\to W^{+}W^{-}$ for t-channel (3) and u-channel (3).
@@
$\mathcal{A}^{2}\simeq\frac{8g^{4}}{M_{W}^{4}}\begin{pmatrix}M_{W}^{4}&\text{tr}^{2}(\mathbb{M})\cdot m_{e\mu}^{2}\\ \text{tr}^{2}(\mathbb{M})\cdot m_{e\mu}^{2}&M_{W}^{4}\end{pmatrix},$ (3.4)
@@
The differential cross-section for this process is:
@@
$\frac{d\sigma}{d\Omega}=\frac{\mathcal{A}^{2}}{64\pi^{2}s}\simeq\frac{g^{4}}{16\pi^{2}}\frac{1}{M_{W}^{6}}\begin{pmatrix}M_{W}^{4}&\text{tr}^{2}(\mathbb{M})\cdot m_{e\mu}^{2}\\ \text{tr}^{2}(\mathbb{M})\cdot m_{e\mu}^{2}&M_{W}^{4}\end{pmatrix}.$ (3.5)
The squared amplitude (3.4) reveals the following:
  • Diagonal terms ($\alpha=\beta$): Suppressed by $M_{W}^{2}$, yielding cross-sections $\propto 10^{-26}$ b/sr.
  • Off-diagonal terms ($\alpha\neq\beta$): Suppressed by $\text{tr}^{2}(\mathbb{M})m_{e\mu}^{2}/M_{W}^{6}\sim 10^{-79}$ b/sr due to tiny neutrino masses.
$\frac{d\sigma_{\mu^{+}e^{-}\to W^{+}W^{-}}}{d\sigma_{e^{+}e^{-}\to W^{+}W^{-}}}\simeq\cfrac{\text{tr}^{2}(\mathbb{M})m_{e\mu}^{2}}{M_{W}^{4}}\simeq 3\cdot 10^{-53}$ (3.6)
@@
The obtained cross section is 53 orders of magnitude lower than that of the reference channel while being flavor-conserving, which precludes the possibility of observing such processes in current and future experiments. This result thus validates neglecting the contributions from flavor-changing neutrino exchange interactions in the analysis of $W$-boson production processes in lepton collisions.
3.2 Neutrino Oscillation
Neutrino oscillations significantly affect processes over large space intervals $\mathbf{x}-\mathbf{y}\gg 0$ and positive time difference $x_{0}-y_{0}>0$.

### v2

@@ Abstract
In neutrino physics, while massive states define neutrino masses, the flavor states participating in weak interactions are governed by an off-diagonal mass matrix. This work examines the complete form of this mass matrix, both for a two-flavor toy model and for the general three-flavor case under two distinct mass hierarchies. Using the Monte Carlo method, we estimate the mass matrix parameters and demonstrate how its structure governs the dependence of the interaction cross section on the mass hierarchy (normal vs inverted). This formalism enables the treatment of processes involving neutrino exchange through a non-diagonal propagator, corresponding to a quantum field theory description. Numerical estimates for a local charged-lepton interaction via virtual neutrino exchange yield a ratio of electron-antimuon (lepton flavor violating) to electron-positron cross sections on the order of $10^{-51}$. Furthermore, the cross section exhibits a fundamental dependence on the lightest neutrino mass, which differs drastically between the two hierarchies. For macroscopic processes, this propagator formalism reproduces the standard neutrino oscillation probability by operating directly with the non-diagonal mass matrix, thereby circumventing the wave-packet formalism and confirming the validity of this approach.
@@ Introduction
Typically, the neutrino which is propagating for the long distances, is considered as a real particle and thus must exist in a mass eigen state. This is achieved through the diagonalization of the mass matrix entering the respective Lagrangian. However, in modern approaches, particularly in Naumov’s [1] and Grimus’ [2], neutrino emerges as a virtual particle. This enables computing these propagators without diagonalizing the mass matrix.
This work investigates the neutrino mass matrix. Analytically, we derive an explicit expression for the mass matrix within the Chau-Keung parametrization of the PMNS matrix, obtaining it as a function of all fundamental parameters (Euler angles $\theta_{ij}$, CP-violating phase $\delta_{\text{CP}}$, and neutrino masses $m_{i}$).
Numerical estimates of the mass matrix were obtained using the Monte Carlo method, incorporating data from the global fit NuFIT 6.0 [3] and Joint neutrino oscillation analysis from T2K and NOvA experiments published in [4]. The study of the mass matrix dependence reveals indications of fine-tuning, which could serve as a criterion for selecting a specific neutrino mass hierarchy.
Furthermore, the work extends the Quantum Field Theory (QFT) approach to lepton number violating processes through an analysis of a two-flavor toy model of Dirac neutrinos. Our complementary approach utilizes a non-diagonal matrix propagator for neutrino flavor states, which serves to validate the formalism based on the non-diagonal mass matrix. A similar methodology was employed by M. Dvornikov [5] in analyzing neutrino oscillations in matter, where the effective mass matrix becomes non-diagonal due to interactions with matter.
We have demonstrated that local processes involving neutrino exchange, such as $l^{-}_{\alpha}l^{+}_{\beta}\to W^{+}W^{-}$, exhibit oscillation-sensitive contributions that are strongly suppressed. The degree of suppression depends critically on the flavors of the initial leptons. For the same-flavor case ($l^{-}_{\alpha}l^{+}_{\alpha}$), the characteristic scale of the cross section is $\mathcal{O}(10^{-11})$ fbarn, which is 11 orders smaller than the leading contribution from direct lepton annihilation obtained by Drutskoy et al. [6]. For different flavors ($l^{-}_{\alpha}l^{+}_{\beta}$), the suppression is far more severe, with the cross section falling to $\mathcal{O}(10^{-62})$ fbarn, effectively rendering it unobservable. The same lepton flavor-violating process was discussed in early studies; for example, the Singhal et al. [7] paper reports a difference of 59 orders of magnitude between cross sections for Majorana neutrinos (which include heavy neutrino states) and our Dirac neutrino approach.
For macroscopic processes, we also consider a process with a change in the flavor of the initial and final leptons via neutrino exchange; however, we employ a non-diagonal propagator. Our result for the amplitude in macroscopic diagrams with two-flavor transitions agrees with that of Kobzarev et al. [8], despite their calculation having used diagonal propagators for massive neutrinos. Both approaches yield an amplitude that reproduces the canonical oscillation dependence
$P_{\text{osc}}\propto\sin^{2}\left(\frac{\Delta m^{2}L}{4E_{\text{avg}}}\right),$
where the oscillation phase is directly linked to the neutrino masses. This formalism bridges the foundations of QFT with experimental interpretation while simultaneously confirming consistency with established quantum mechanical (QM) results.
@@ 1 Dirac neutrino propagator
The propagator in momentum space, analogous to Krivoruchenko and Šimkovic’s in [9], is defined as $\mathbb{K}^{-1}(p)\equiv\mathbb{D}(p)$ :
@@ 1.1 Two-flavor propagator
We consider the specific case of interaction between electron flavor ($\alpha=e$) and muon flavor ($\alpha=\mu$), with a real symmetric mass matrix:
@@ 2 Three flavor mass matrix
Analyzing the propagator for Dirac neutrino with three flavors in its off-diagonal form presents several complications. However, it is useful to trace the connection between the mass matrix and the mixing angles for three flavors. When mass matrix have only real elements, we have 6 independent variables that can be expressed through 6 parameters: 3 masses neutrino and 3 mixing angles. But we have a Hermitian mass matrix with nine different variables and (2.2) is expressed through the only seven independent parameters: 3 masses, 3 mixing angles and CP-phase $\delta$. We expect there are two relation between mass matrix elements. In the general case, the mass matrix is Hermitian:
@@
2.1 Mass matrix in Chau-Keung parametrization
Standard Chau-Keung parametrization[10] of Pontecorvo-Maki-Nakagawa-Sakata (PMNS) matrix $V_{\text{PMNS}}$ has three Euler angles $\theta_{12},\theta_{23},\theta_{13}$ and Dirac CP-phase $\delta$:
@@ 2.2 Numerical estimates
Neutrino mass measurements do not yield absolute mass values but squared mass differences. The absolute mass scale is set by the lightest neutrino mass in a given mass ordering.
Figure 1: analogue of the Mandelstam plane
For numerical estimates, we employ the standard Chau-Keung parametrization[10], where the mixing angles and the CP-violating phase are experimentally measured quantities. The current best-fit values for these parameters are provided by the global analysis of NuFIT v6.0[3], as well as by a recent Joint neutrino oscillation analysis from the T2K and NOvA experiments published in Nature[4]. We have 2 different situation for Inverted and Normal Ordering mass states.
@@
$\theta_{12}\,/^{\circ}$ $3.68^{+0.73}_{-0.70} 33$$3.68^{+0.73}_{-0.70}$
$\theta_{23}\,/^{\circ}$ $48.4^{+1.7}_{-2.8}$ $48.4^{+1.7}_{-2.8}$
@@
$\cfrac{\Delta m^{2}_{31}}{10^{-3}\,\text{eV}^{2}}$ $2.513^{+0.021}_{-0.019}$ $2.409^{+0.020}_{-0.020}$
$\cfrac{\Delta m^{2}_{32}}{10^{-3}\,\text{eV}^{2}}$ $2.438^{+0.021}_{-0.019}$ $2.484^{+0.020}_{-0.020}$
Table 1: Parameters from NuFIT 6.0 (2024)[3]
and Joint neutrino oscillation analysis (2025)[4].
The mass matrix values and their uncertainties are estimated using a Monte Carlo technique. First, probability distributions for the mixing angles and mass-squared differences are constructed, accounting for the mass hierarchy—specifically, the dependence on the lightest neutrino mass, which is bounded by the cosmological upper limit. These distributions are then used to generate mass matrices as functions of the lightest neutrino mass.
The resulting distribution of the mass matrix elements yields their $1\sigma$ uncertainties. The dependence of neutrino mass matrix elements on the lightest neutrino mass for both normal and inverted hierarchies is shown in Appendix A.
Normal hierarchy mass matrix ($\mathbb{M}_{\text{NH}}\times 10^{2}\text{ eV}$), defined under the assumption $m_{1}=0$
[…]
Ratio for the inverse hierarchy ($\mathbb{M}_{\text{IH}}\times 10^{2}\text{ eV}$) with respect to the $m_{3}=0$ case:
[…]
@@
$\begin{split}&m_{3}=0\times 10^{-2}\text{ eV}\ \sim\ \Re(m_{e\mu})=-4.07\cdot 10^{-5},\\ &m_{3}=2\times 10^{-2}\text{ eV}\ \sim\ \Re(m_{e\mu})=3.66\cdot 10^{-5},\\ &m_{3}=4\times 10^{-2}\text{ eV}\ \sim\ \Re(m_{e\mu})=5.5\cdot 10^{-5}.\end{split}$ (2.8)
@@
Figure 2: Plots of $\Re(m_{e\mu})$ for both neutrino mass hierarchies: normal hierarchy on the left and inverted hierarchy on the right, as a function of the lightest neutrino mass.
@@
All non-diagonal elements of neutrino mass matrix exhibit power-law decay towards zero without sign inversion. However, $\Re(m_{e\mu})$ displays distinct behavior in both mass hierarchies. For the Normal Hierarchy (NH), we observe an initial local minimum followed by asymptotic convergence to zero. In the Inverted Hierarchy (IH), $\Re(m_{e\mu})$ undergoes sign reversal, reaches a global maximum beyond cosmological bounds on $m_{3}$, and subsequently approaches zero from above. This pronounced peak, while cosmologically nonphysical, presents a striking feature. The inverse hierarchy is related to some kind of fine-tuning for matrix. This may be an additional argument in choosing the normal hierarchy (NH).
@@ 3 Weak Interaction Process
where $\mathcal{P}_{L}=(1-\gamma_{5})/2$ is the left-handed projector, satisfying $\mathcal{P}_{L}\mathcal{P}_{R}=0$. In Feynman diagrams, vertices include $\gamma^{\mu}\mathcal{P}_{L}$. Consequently, in the amplitude squared, terms with constants between vertices vanish due to:
@@ 3.1 Process $l_{\alpha}^{-}l_{\beta}^{+}\to W^{+}W^{-}$
We consider crossing of two charged lepton and antilepton ($e^{\mp}$, $\mu^{\pm}$) scattering into two W-bosons and evaluate the contribution of neutrino oscillations to this local process.
Figure 3: Diagram of $l_{\alpha}^{-}l_{\beta}^{+}\to W^{+}W^{-}$ for t-channel.
@@
$\mathcal{A}^{2}\simeq\frac{8g^{4}}{M_{W}^{4}}\begin{pmatrix}M_{W}^{4}&\text{tr}^{2}(\mathbb{M})\cdot|m_{e\mu}|^{2}\\ \text{tr}^{2}(\mathbb{M})\cdot|m_{e\mu}|^{2}&M_{W}^{4}\end{pmatrix}$, (3.4)
@@
The differential cross section for this process is:
@@
$\frac{d\sigma}{d\Omega}=\frac{\mathcal{A}^{2}}{64\pi^{2}s}\simeq\frac{g^{4}}{16\pi^{2}}\frac{1}{M_{W}^{6}}\begin{pmatrix}M_{W}^{4}&\text{tr}^{2}(\mathbb{M})\cdot|m_{e\mu}|^{2}\\ \text{tr}^{2}(\mathbb{M})\cdot|m_{e\mu}|^{2}&M_{W}^{4}\end{pmatrix}$. (3.5)
Considering the numerical estimates for $|m_{e\mu}|^{2}$ from (2.6) and (2.7), the amplitude squared (3.4) will yield: For the case of identical flavors (diagonal elements), the cross sections are suppressed by $M_{W}^{2}$ for both hierarchies and are approximately equal to $d\sigma\simeq 10^{-11}$ fbarn.
Regarding the off-diagonal element, the cross sections are suppressed by more than $M_{W}^{6}$ and are approximately equal to $d\sigma\simeq 10^{-62}$ fbarn for both hierarchies.
Numerical estimates of the cross section dependence on the increasing lightest neutrino mass show that, on the one hand, the approximate cross section increases with $m_{1}$ for the normal hierarchy (fig. 4). On the other hand, for the inverted hierarchy, this value decreases as $m_{3}$ approaches its maximum value (fig. 4).
Figure 4: Plots of $\frac{d\sigma}{d\Omega}$ for both neutrino mass hierarchies: normal (left) and inverted (right), as a function of the lightest neutrino mass $m_{1}$ and $m_{3}$, respectively.
It should be noted that the electron-positron annihilation process with W-boson pair production was considered in [6], where the obtained cross section is 11 orders of magnitude larger. This discrepancy arises primarily because we do not consider heavy neutrinos, and secondly, because we only examine neutrino exchange, which is significantly smaller than the cross sections for W-boson pair annihilation and their interference.
As for processes with lepton number violation, a similar reaction was studied in [7] for Majorana neutrinos. For Dirac neutrinos, the difference in cross sections amounts to 59 orders of magnitude, which is also attributed to the consideration of heavy neutrinos.
The ratio of flavor-changing cross sections to the electron-positron cross section for the two mass hierarchies
$\frac{d\sigma_{\mu^{+}e^{-}\to W^{+}W^{-}}}{d\sigma_{e^{+}e^{-}\to W^{+}W^{-}}}\simeq\cfrac{\text{tr}^{2}(\mathbb{M})m_{e\mu}^{2}}{M_{W}^{4}}$ (3.6)
@@
corresponds to value on the order of $10^{-51}$ for both hierarchy. Such a small value of the cross section ratio excludes the possibility of observing these processes in current and future experiments. This result thus validates neglecting the contributions from flavor-changing neutrino exchange interactions in the analysis of $W$-boson production processes in lepton collisions.
3.2 Neutrino Oscillations
The manifestation of lepton flavor-violating processes via virtual neutrino exchange appears in the s-channel of macro-diagram (fig. 5), where the spatial $\mathbf{x}-\mathbf{y}\gg 0$ and $x_{0}-y_{0}>0$ separation between the source $S$ and the detector $D$ plays a significant role in neutrino oscillations.
Figure 5: Diagram of $l_{\alpha}W^{+}\to l_{\beta}W^{+}$ for s-channel.
For simplicity, we again turn to a toy model with two flavors, which allows us to work quite straightforwardly with matrix forms. The neutrino propagator in such a process must be considered in coordinate space, which explicitly highlights the distance L between the detector and the source.
$\mathbb{D}(x-y)=\frac{1}{4\pi^{2}\text{L}}\sum_{k=1,2}{(-1)^{k}}\frac{\text{p}_{k}}{2\omega_{k}}\left[\gamma^{0}\omega_{k}+\mathbb{M}-\left(\text{p}_{k}+\cfrac{i}{\text{L}}\right)\mathbf{n}_{\text{L}}\bm{\gamma}\right]\mathbb{A}_{k}\ e^{i\text{p}_{k}\text{L}}$ (3.7)
where the momenta $p_{k}$ correspond to the $k$ neutrino masses
$\text{p}_{k}=\sqrt{E_{\text{avg}}^{2}-m_{k}^{2}}$ (3.8)
and $E_{\text{avg}}$ denote the average energy at the detector and source, respectively. The matrices $\mathbb{A}_{k}$ are expressed through the mixing angle $\theta$, defined below in (3.13)
$\mathbb{A}_{1}\equiv\begin{pmatrix}-\cos^{2}\theta&\cfrac{\sin(2\theta)}{2}\\ \cfrac{\sin(2\theta)}{2}&-\sin^{2}\theta\\ \end{pmatrix},\quad\mathbb{A}_{2}\equiv\begin{pmatrix}\sin^{2}\theta&\cfrac{\sin(2\theta)}{2}\\ \cfrac{\sin(2\theta)}{2}&\cos^{2}\theta\\ \end{pmatrix}$. (3.9)
A propagator of similar form to (3.7) was derived by Kobzarev et al.  [^1] See formula (1.5) of Appendix 1 in [8] in a general diagonalized representation for three flavors:
$\mathbb{D}(x-y)\propto\frac{1}{4\pi\text{L}}\left[\varepsilon\gamma_{0}+m_{A}-\left(\text{p}_{A}+\cfrac{i}{\text{L}}\right)\mathbf{n}_{\text{L}}\bm{\gamma}\right]U_{iA}^{\dagger}U_{kA}e^{i\text{p}_{A}\text{L}}.$ (3.10)
The specific form of the matrix propagator, consisting of a sum, leads to an amplitude of the form:
$\mathcal{A}_{\alpha\beta}\propto\frac{1}{\text{L}}\left(\mathbb{A}_{2}^{\alpha\beta}\exp\left\{i\text{L}\sqrt{E_{\text{avg}}^{2}-m_{2}^{2}}\right\}-\mathbb{A}_{1}^{\alpha\beta}\exp\left\{i\text{L}\sqrt{E_{\text{avg}}^{2}-m_{1}^{2}}\right\}\right),$ (3.11)
the square of which produces interfering terms responsible for flavor change in the process.